OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 133, 126…, if there are 33 terms?

OpenStudy (anonymous):

d = 7

OpenStudy (anonymous):

231 i think would be the sum

OpenStudy (anonymous):

s=n(a1+an)/2 let n = 33 an = 140+(32)-7= -84 s=33(140-84)/2=924

OpenStudy (anonymous):

If you substitute the explicit definition into Gauss' formula you get a formula that does not require the last term\[s=\frac{n}{2}(2a _{1}+(n-1)d)\]\[s=\frac{33}{2}\left( 2*140+(33-1)(-7) \right)=924\]